a rock is thrown horizontally from the top of a cliff 88 m high, with a horizontal speed of 25 m/s. with what velocity does the rock hit?

1
Expert's answer
2022-05-11T13:52:41-0400
"v=\sqrt{v_0^2+2gh}\\
=\sqrt{25^2+2*9.8*88}=48\:\rm m/s"